y=(u-2)3 and u=1/2x=1


1
Expert's answer
2021-07-03T08:05:47-0400

"y=(u-2)^3"

"u=\frac{1}{2}x-1"

"y=(\frac{1}{2}x-1-2)^3"

"y=(\frac{1}{2}x-3)^3"

But x=1

"y=(\frac{1}{2}-3)^3"

"y=-\frac{125}{8}"